Skyler invests $4700 in two different accounts. The first account paid 5 %, the second account paid 6 % in interest. At the end of the first year he had earned $271 in interest. How much was in each account?
5%-
6%-

Answer by Boreal(15235)

x at 5%
4700-x at 6%
.05x+.06(4700-x)=271
.05x-.06x+282=271
-.01x=-11
x=$1100 @ 5% or $55
4700-x=3600@6% or $216
5%:$1100
6%:$3600